# Source reconstruction algorithms for coupled parabolic systems from internal measurements of one scalar state

### v3

@@ 5. Numerical results
In this section, we present numerical tests related to Problems 1 and 2 proved in Theorems 1 and 3, respectively. Since our theoretical results proved in previous sections are in $\mathbb{R}^{n}$, with $n>1$ when the coupling is given for a constant matrix $Q$, and $n=1$ for the case where the coupling has space-dependent coefficients on the matrix $Q$, we perform 1D and 2D numerical experiments under different configurations.
In concordance with Problems 1 and 2, we recover the source term $F(x)=(f_{1}(x),f_{2}(x))^{t}$ as the solution of the following minimization problem:
@@
where $Y_{\text{obs}}:=(y_{1,\text{obs}},y_{2,\text{obs}})^{t}$ are observation data on the subdomain $\mathcal{O}\times(0,T)$ (or also so-called partial measurements of the state $Y=(y_{1},y_{2})^{t}$), $Q$ is the coupling matrix given by
$Q=\begin{pmatrix}0&q_{12}\\ q_{21}&0\end{pmatrix}$,
with $q_{12}\neq 0,q_{21}\neq 0$, and $\sigma(t):=1+\frac{1}{2}\cos\left(\frac{4\pi t}{T-t_{0}}\right)$ for $t<T-t_{0}$ is the sinusoidal function selected for all experiments (see [28, 27]). The functional $J$ is selected in this form based on the theoretical requirements for optimization problems and following ideas from [27]. Additionally, the Lipschitz-type stability result presented in Appendix B allows us to clarify the structure of the function $J(Y,F)$ defined in (5.1). The PDE constraining the above minimization problem (5.2) is solved using finite elements with conforming piecewise linear Lagrange finite element basis functions in 1D and 2D domains. The finite element schemes are solved using the FEniCS library [4]. For minimizing the functional (5.1), we employed a steepest descent method [12, 36], with a penalty term $k=10^{5}$ and a fixed step size $10^{-4}$.
@@ 5.1. Source reconstruction in 1D
Figures 2 and 3 show the numerical results of source reconstructions considering two different observation domains, namely $\mathcal{O}_{1}=(0.5,0.9)$ and $\mathcal{O}_{2}=(0.2,0.4)\cup(0.6,0.8)$, and two distinct coupling scenarios. Figure 2 illustrates reconstructions obtained using a coupling matrix $Q$ defined by cubic polynomial terms, with $q_{12}(x)=x(x-1)(5x-2)+0.5$ and $q_{21}(x)=x(1-x)(5x-3)+0.5$. In this setting, reconstructions using measurements from both components consistently exhibit robust performance and relatively low reconstruction errors. On the other hand, reconstructions based on single-component measurements appear more sensitive to both the location and size of the observation subdomain, leading in some cases to larger reconstruction errors, particularly outside the observation region.
Similarly, Figure 3 presents reconstructions using a linear coupling configuration with $q_{12}(x)=4x-2$ and $q_{21}(x)=-4x+2$. These reconstructions exhibit trends analogous to those observed in the cubic polynomial case, though with slightly improved accuracy overall. Moreover, it can be observed from Figures 2 and 3 that reconstruction quality significantly improves when the observation subdomain covers regions with stronger variations of the source function. This indicates that source reconstruction accuracy is linked not only to the size and location of the observation domain but also to its alignment with the spatial structure of the source.
(a) rel. error $12,7\%$
(b) rel. error $44,6\%$
(c) rel. error $45,7\%$
(d) rel. error $9,7\%$
(e) rel. error $20,9\%$
(f) rel. error $60,9\%$
(g) rel. error $2,9\%$
(h) rel. error $24,9\%$
(i) rel. error $42,9\%$
@@
(k) rel. error $21,4\%$
(l) rel. error $44,9\%$
Figure 2. Source reconstructions using coupling terms $q_{12}(x)=x(x-1)(5x-2)+0.5$ and $q_{21}(x)=x(1-x)(5x-3)+0.5$. First column is the reconstruction of $F_{1}(x)$ and $F_{2}(x)$ observing both components, i.e., $(y_{1,\text{obs}},y_{2,\text{obs}})^{t}$. The second column shows the reconstruction of $F_{1}(x)$ and $F_{2}(x)$ observing the first component $y_{1,\text{obs}}$, and the last one is the reconstruction observing the second component $y_{2,\text{obs}}$ only. The highlighted gray areas correspond to the observation domain $\mathcal{O}$.
@@
Figure 3. Source reconstructions using coupling terms $q_{1}(x)=4x-2$ and $q_{2}(x)=-4x+2$. First column is the reconstruction of $F_{1}(x)$ and $F_{2}(x)$ observing both components, i.e., $(y_{1,\text{obs}},y_{2,\text{obs}})^{t}$. The second column shows the reconstruction of $F_{1}(x)$ and $F_{2}(x)$ observing the first component $y_{1,\text{obs}}$, and the last one is the reconstruction observing the second component $y_{2,\text{obs}}$ only. The highlighted gray areas correspond to the observation domain $\mathcal{O}$.
@@ 5.2. Source reconstruction varying the regularization term
We analyze the effect of the regularization term $k$ on the reconstruction quality in a one-dimensional setting. Specifically, we consider the sinusoidal source $F_{1}(x):=(\sin(2\pi x),-\sin(2\pi x))^{t}$, the observation domain $\mathcal{O}_{1}:=(0.5,0.9)$, and use cubic polynomial coupling terms as previously presented in Figure 2. All other parameters remain consistent with those from earlier experiments. Our goal here is to investigate how different choices of the regularization parameter $k$ influence the quality of the reconstructed source, particularly in regions inside and outside the observation subdomain.
@@
Figure 4 displays reconstruction results using measurements from both state components, with the regularization parameter $k$ varying across the set $\{10^{2},10^{3},10^{4},10^{5}\}$. As illustrated, increasing the value of $k$ generally leads to significant improvements in reconstruction quality. More precisely, larger $k$ values yield more accurate reconstructions, reducing both amplitude discrepancies and oscillatory artifacts, notably outside the observation region. However, an excessively small regularization parameter (e.g., $k=10^{2}$) results in pronounced reconstruction errors, with oscillations and deviations becoming particularly evident far from the observation domain. Thus, selecting an adequately large regularization parameter is essential for balancing accuracy and stability in source reconstruction tasks.
(a) $k=10^{2}$; rel. error $68,8\%$
(b) $k=10^{3}$; rel. error $53,1\%$
(c) $k=10^{4}$; rel. error $11,3\%$
(d) $k=10^{5}$; rel. error $12,7\%$
Figure 4. Source reconstruction for $F_{1}(x)$ under different values of the parameter $k$ in the functional (5.1). The highlighted gray areas correspond to the observation subdomain $\mathcal{O}$.
5.3. Source reconstruction under noisy observations
Measurements in practical applications typically contain some level of noise resulting from, e.g., instrumental limitations, environmental interference, or data acquisition processes. To evaluate the robustness of our reconstruction method under realistic conditions, we analyze the impact of different noise levels on the reconstruction quality. Specifically, we introduce Gaussian noise into the finite element approximations of the observations. At each time step $j\Delta t$ the finite element solution is given by $y_{h}(j\Delta t):=\sum_{i=1}^{m}c^{j}_{i}\phi_{i}^{\text{FE}}$, where $\phi_{i}^{\text{FE}}$ represents the $i$-th basis function of the finite element space. To simulate noisy measurements, the coefficients $c_{i}^{j}$ are modified as follows:
$c_{noisy,i}^{j}=c_{i}^{j}+r*\textup{std}({\bf c})*\zeta_{i}^{j},$
where ${\bf c}=(c_{i}^{j})_{i,j}$, $r$ controls the noise intensity, std represents the standard deviation, and $\zeta_{i}^{j}$ are random variables following a normal standard distribution, independently generated for each coefficient and each time step. Consequently, the synthetic observations $Y_{\text{obs}}:=(y_{1,\text{obs}},y_{2,\text{obs}})^{t}$ used in the minimization problem (5.1) are computed from these noise-corrupted finite element solutions.
Figure 5 illustrates how different levels of noise affect the quality of the source reconstruction, comparing results obtained from observations of both state components versus observations of a single component. As expected, reconstructions from noise-free measurements using both components achieve low reconstruction errors. However, reconstructions based on both components exhibit greater sensitivity to noise, particularly within the observation region. In contrast, single-component reconstructions show lower sensitivity to noise, resulting in a more robust strategy under noisy observations.
In addition, Figure 5 highlights that moderate noise levels (e.g., $5\%$ or $10\%$) introduce noticeable reconstruction errors within the observation domain for two-component reconstructions, suggesting a need for careful noise management or additional regularization in practical scenarios. At higher noise levels (e.g., $20\%$), errors become pronounced even for single-component reconstructions, particularly affecting the accuracy outside the observation subdomain.
(a) rel. error $2,7\%$ (no noise)
(b) rel. error $21,4\%$ (no noise)
(c) rel. error $21,4\%$ (no noise)
(d) rel. error $11,4\%$ ($5\%$ noise)
(e) rel. error $21,0\%$ ($5\%$ noise)
(f) rel. error $20,5\%$ ($5\%$ noise)
(g) rel. error $21,9\%$ ($10\%$ noise)
(h) rel. error $23,5\%$ ($10\%$ noise)
(i) rel. error $22,4\%$ ($10\%$ noise)
(j) rel. error $43,1\%$ ($20\%$ noise)
(k) rel. error $34,1\%$ ($20\%$ noise)
(l) rel. error $32,3\%$ ($20\%$ noise)
Figure 5. Source reconstruction for $F_{1}(x)$ using linear coupling terms $q_{1}(x)=4x-2$ and $q_{2}(x)=2-4x$ under different levels of noise: $0\%$, $5\%$, $10\%$, and $20\%$. The first column shows the reconstruction observing both components. The second column shows the reconstruction observing the first component, and the last one displays the reconstruction observing the second component. The highlighted area corresponds to the observation domain $\mathcal{O}$.
@@ 5.4. Source reconstruction in 2D
In this subsection, we analyze two-dimensional source reconstructions on the unit square domain $\Omega=(0,1)^{2}$, using parameters $T=0.5$, and $\nu=0.1$. For the numerical discretization, we utilize a regular triangular mesh composed of $3200$ elements, with a temporal discretization step size $\Delta t=0.01$. We test two observation domains: $\mathcal{O}_{1}=(0.5,0.9)\times(0.1,0.9)$ and $\mathcal{O}_{2}=((0.2,0.4)\cup(0.6,0.8))\times(0.3,0.7)$. The reconstruction accuracy is evaluated using the following two source functions:
  1. a)
    $F_{3}(x,y):=(\sin(2\pi x)\sin(2\pi y),-\sin(2\pi x)\sin(2\pi y))^{t}$.
  2. b)
    $F_{4}(x,y):=(f_{4,1}(x,y),f_{4,2}(x,y))^{t}$ defined by:
[…]
Figures 6 and 7 illustrate the numerical reconstruction results obtained for these sources. Specifically, Figure 6 shows reconstructions of the sinusoidal source $F_{3}$ using constant coupling terms $q_{12}(x,y)=4$ and $q_{21}(x,y)=2$. The two top rows present results based on observations within the domain $\mathcal{O}_{1}$, while the two bottom rows show reconstructions derived from measurements in the domain $\mathcal{O}_{2}$. Similarly, Figure 7 displays reconstructions of the discontinuous source $F_{4}$ using symmetric coupling terms $q_{12}(x,y)=q_{21}(x,y)=3$.
@@
Figure 6. Source reconstruction for $F_{3}(x,y)$ using coupling terms $q_{12}(x,y)=4$ and $q_{21}(x,y)=2$. The first two rows show reconstructions with measurements on $\mathcal{O}_{1}$ (top), and the last two rows show reconstructions with measurements on $\mathcal{O}_{2}$ (bottom). The first column displays the sources and the location of the observation domains in a white dashed line. The second, third, and fourth columns display the reconstructions with both components, followed by the reconstructions with the first and second components, respectively.
Figure 7. Source reconstruction for $F_{4}(x,y)$ using coupling terms $q_{12}(x,y)=3$ and $q_{21}=3$. The first two rows show reconstructions with measurements on $\mathcal{O}_{1}$ (top), and the last two rows show reconstructions with measurements on $\mathcal{O}_{2}$ (bottom). The first column displays the sources and the location of the observation domains in a white dashed line. The second, third, and fourth columns display the reconstructions with both components, followed by the reconstructions with the first and second components, respectively.
@@ 6. Conclusions
In this paper, we introduced and analyzed inverse problems aimed at recovering spatial source distributions in coupled parabolic systems. We addressed two types of coupling: one involving a potential matrix with either constant or spatially dependent entries, and the other involving coupling through the main operator. A significant contribution of our work is the development of source reconstruction algorithms based on local internal measurements from one scalar component of the system’s state. To the best of our knowledge, there are no previous studies on reconstruction algorithms for these sources.
One of the primary advantages of our approach is the minimal requirement for the measured state variables needed to reconstruct the source coefficients. Nevertheless, several related theoretical questions remain unresolved, particularly those concerning stability and uniqueness of the inverse source problem using the methodology via spectral analysis, control null results, and Volterra equations. This lack of resolution is not unexpected; even in the scalar case analyzed in [28], achieving stability required stringent assumptions on the temporal derivative $\sigma^{\prime}(t)$. Specifically, a stability result was obtained when $\sigma^{\prime}(t)$ decreases exponentially in $(0,T)$, i.e., $\sigma^{\prime}(t)\approx e^{-1/t}$ (see [28, Proof of Theorem 1.3, Step 4]). This particular choice of $\sigma^{\prime}(t)$ is closely linked to the null controllability problem since the control function cost has an exponential dependence of the form $e^{1/T}$ (see [28, Lemma 1.1]). Consequently, establishing stability via our methodology would similarly require precise control-cost estimates for coupled parabolic systems. Moreover, such estimates must simultaneously allow appropriate bounds on $\sigma(t)$ and $\sigma^{\prime}(t)$ while ensuring $\sigma(T)\neq 0$. Currently, however, obtaining these estimates remains an open problem, as the control function cost involves an exponential term of the form $e^{T+1/T}$. To provide a positive answer regarding the stability property, we have included a Lipschitz-type stability result for Problem 1 via Carleman inequalities in the Appendix B.

### v4

@@ 5. Numerical results
In this section, we present numerical tests related to Problems 1 and 2 proved in Theorems 1 and 2, respectively. Since our theoretical results proved in previous sections are in $\mathbb{R}^{n}$, with $n>1$ when the coupling is given for a constant matrix $Q$ and $n=1$ for the case where the coupling has space-dependent coefficients on the matrix $Q$, we perform 1D and 2D numerical experiments under different configurations.
The following section presents an optimization-based methodology designed for the complete numerical analysis, in contrast to the framework discussed in the previous section, which is grounded in spectral analysis, the null controllability property, and Volterra equations. Further details on this approach are given in subsection  5.4. In concordance with Problems 1 and 2, we recover the source term $F(x)=(f_{1}(x),f_{2}(x))^{t}$ as the solution of the following minimization problem:
@@
where $Y_{\text{obs}}:=(y_{1,\text{obs}},y_{2,\text{obs}})^{t}$ are observation data on the subdomain $\mathcal{O}\times(0,T)$ (or also so-called partial measurements of the state $Y=(y_{1},y_{2})^{t}$). We also define
$Q=\begin{pmatrix}q_{11}&q_{12}\\ q_{21}&q_{22}\end{pmatrix},\quad\text{and}\quad\sigma(t):=\begin{cases}1+\frac{1}{2}\cos\left(\frac{4\pi t}{T-t_{0}}\right)&t<T-t_{0}\\ \frac{3}{2}&t\geq T-t_{0}\end{cases},$
where $Q$ is the coupling matrix and $\sigma$ is the function selected for all experiments (see [25, 24]). The functional $J$ is selected in this form based on the theoretical requirements for optimization problems and following ideas from [24]. Additionally, the Lipschitz-type stability result presented in Appendix A allows us to clarify the structure of the function $J(Y,F)$ defined in (5.1). The PDE constraining the above minimization problem (5.2) is solved using finite elements with conforming piecewise linear Lagrange finite element basis functions in 1D and 2D domains. The finite element schemes are solved using the FEniCS library [4]. For minimizing the functional (5.1), we employed a steepest descent method [11, 33], with a penalty term $k=10^{5}$ and a fixed step size $10^{-4}$.
@@ 5.1. Source reconstruction in 1D
Figure 2 illustrates reconstructions obtained using a coupling matrix defined by $q_{11}(x)=q_{12}(x)=q_{22}(x)=0$ and $q_{21}(x)=-x^{3}+4x^{2}-3x+1$ and and two different observation domains, namely $\mathcal{O}_{1}=(0.5,0.9)$ and $\mathcal{O}_{2}=(0.2,0.4)\cup(0.6,0.8)$. In this setting, reconstructions using measurements from both components consistently exhibit robust performance and relatively low reconstruction errors. On the other hand, reconstructions based on single-component measurements appear more sensitive to both the location and size of the observation subdomain, leading in some cases to larger reconstruction errors, particularly outside the observation region. Note that, in the second column, the reconstruction of the source term $f_{2}$ from the first state does not yield a satisfactory error. This suggests that the coupling term and the observation domain are not sufficiently strong to achieve an efficient reconstruction of the second source from the first state. This observation, arising from the numerical experiments, also highlights open problems in the literature on inverse problems, namely, questions regarding the optimal location of the observation subdomain and the possibility of determining an optimal matrix configuration for the coupling.
Figure 3 presents reconstructions using a different coupling configuration $Q(x)$ given by $q_{11}(x)=q_{22}(x)=0$, $q_{12}(x)=4x-2$, and $q_{21}(x)=-4x+2$. Note that this choice for the coupling matrix describes a setting more general than presented in Lemma 1. In this setting, reconstructions from measurements in both components consistently exhibit robust performance and relatively low reconstruction errors. On the other hand, reconstructions based on single-component measurements appear more sensitive to both the location and size of the observation subdomain, leading in some cases to larger reconstruction errors, particularly outside the observation region Moreover, it can be observed from Figures 2 and 3 that reconstruction quality significantly improves when the observation subdomain covers regions with stronger variations of the source function. This indicates that source reconstruction accuracy is linked not only to the size and location of the observation domain but also to its alignment with the spatial structure of the source.
(a) rel. error $11,7\%$
(b) rel. error $71,3\%$
(c) rel. error $42,3\%$
(d) rel. error $2,5\%$
(e) rel. error $70,7\%$
(f) rel. error $28,3\%$
(g) rel. error $2,5\%$
(h) rel. error $70,7\%$
(i) rel. error $28,3\%$
@@
(k) rel. error $70,7\%$
(l) rel. error $35,8\%$
Figure 2. Source reconstructions when the coupling matrix is given by $q_{11}(x)=q_{12}(x)=q_{22}(x)=0$ and $q_{21}(x)=-x^{3}+4x^{2}-3x+1$. First column is the reconstruction of $F_{1}(x)$ and $F_{2}(x)$ observing both components, i.e., $(y_{1,\text{obs}},y_{2,\text{obs}})^{t}$. The second column shows the reconstruction of $F_{1}(x)$ and $F_{2}(x)$ observing the first component $y_{1,\text{obs}}$, and the last one is the reconstruction observing the second component $y_{2,\text{obs}}$ only. The highlighted gray areas correspond to the observation domain $\mathcal{O}$.
@@
Figure 3. Source reconstructions when the coupling matrix is given by $q_{11}(x)=q_{22}(x)=0$, $q_{12}(x)=4x-2$, and $q_{21}(x)=-4x+2$. . First column is the reconstruction of $F_{1}(x)$ and $F_{2}(x)$ observing both components, i.e., $(y_{1,\text{obs}},y_{2,\text{obs}})^{t}$. The second column shows the reconstruction of $F_{1}(x)$ and $F_{2}(x)$ observing the first component $y_{1,\text{obs}}$, and the last one is the reconstruction observing the second component $y_{2,\text{obs}}$ only. The highlighted gray areas correspond to the observation domain $\mathcal{O}$.
@@ 5.2. Source reconstruction varying the regularization term
We analyze the effect of the regularization term $k$ on the reconstruction quality in a one-dimensional setting. Specifically, we consider $F_{1}(x):=(\sin(2\pi x),-\sin(2\pi x))^{t}$, the observation domain $\mathcal{O}_{1}:=(0.5,0.9)$, and the matrix coupling presented in Figure 3. All other parameters remain consistent with those from earlier experiments. Our goal here is to investigate how different choices of the regularization parameter $k$ influence the quality of the reconstructed source, particularly in regions inside and outside the observation subdomain.
$k$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$
rel. error $67,6\%$ $51,9\%$ $27,2\%$ $12.6\%$ $14,5\%$
Table 1. Error reconstruction for $F_{1}(x)$ under different values of the parameter $k$ for the functional $J$.
@@
Table 1 displays reconstruction results using measurements from both state components, with a regularization parameter $k$ in $\{10^{2},10^{3},10^{4},10^{5},10^{6}\}$. As illustrated, a large $k\sim 10^{5}$ produces significant improvements in the reconstruction. Excessively large $k$ values may lead to overfitting the observations without fully capturing the PDE problem. However, an excessively small regularization parameter (e.g., $k=10^{2}$) results in pronounced reconstruction errors, with oscillations and deviations becoming particularly evident far from the observation domain. Thus, selecting an adequately large regularization parameter is essential for balancing accuracy and stability in source reconstruction tasks.
@@ 5.3. Source reconstruction in 2D
In this subsection, we analyze source reconstructions on the unit square domain $\Omega=(0,1)^{2}$, using parameters $T=0.5$ and $\nu=0.1$. For the numerical discretization, we utilize a regular triangular mesh composed of $3200$ elements, with a temporal discretization step size $\Delta t=0.01$. We test two sets of observation domains: $\mathcal{O}_{1}=(0.5,0.9)\times(0.1,0.9)$ and $\mathcal{O}_{2}=((0.2,0.4)\cup(0.6,0.8))\times(0.3,0.7)$. The reconstruction accuracy is evaluated using the source function $F_{3}(x,y):=(\sin(2\pi x)\sin(2\pi y),-\sin(2\pi x)\sin(2\pi y))^{t}$.
Figures 4 and 5 illustrate the numerical reconstruction results obtained for the above source. Figure 4 shows reconstructions from a constant coupling term given by $q_{11}=1$, $q_{12}=4$, $q_{21}=0$, and $q_{22}=1$. The two top rows present results based on observations within the domain $\mathcal{O}_{3}=(0.3,0.5)\times(0.2,0.8)$, while the two bottom rows show reconstructions derived from measurements in the domain $\mathcal{O}_{4}=((0.2,0.4)\cup(0.6,0.8))\times(0.2,0.8)$. Figure 5 shows reconstructions from a more general constant coupling term using the coupling matrix $q_{11}=q_{22}=0$, $q_{12}(x,y)=4$ and $q_{21}(x,y)=2$, with $\mathcal{O}_{5}=(0.5,0.9)\times(0.1,0.9)$ in the two top rows and $\mathcal{O}_{6}=((0.2,0.4)\cup(0.6,0.8))\times(0.3,0.7)$ in the two bottom rows. Note that the last choice for $Q$ describes a setting more general than presented in Lemma 2.
@@
5.4. Implementation notes
Through this numerical section, we employ a minimization-based formulation that circumvents the need to repeatedly solve large-scale controllability problems and Volterra systems, while still preserving the essential structural concepts of the theoretical framework. This methodological shift is due to the fact that implementing the reconstruction formulas on realistic meshes with practical time resolutions becomes computationally prohibitive. In what follows, we provide a detailed analysis of the computational cost associated with the controllability problems and the Volterra systems:
  1. (1)
    In the null controllability problems, for each $\tau\in(0,T)$ and each eigenfunction indexed by $k$, we must solve a controllability system with initial data $\Psi_{k}^{0}$ (constructed from Laplace eigenfunctions). Implementing this via the Hilbert Uniqueness Method (HUM) requires both time discretization and space discretization. For a system with $n$ equations over a mesh with $M$ nodes, the cost is roughly $O(M\times n\times L)$ per $k$, where $L$ is the number of iterations required for convergence. Note that this grows quickly with mesh refinement or larger $n$.
  2. (2)
    For the Volterra equations, once the control functions $U^{(\tau)}$ are computed, $n$ copies of the Volterra equations (first or second kind, depending on $\sigma(0)$ must be solved for each component and for each $k$. Even with efficient quadrature (e.g., trapezoidal rule), the cost is approximately $O(M\times N\times n\times(\ell+1))$ per $k$, where $N$ is the number of time steps and $\ell$ indexes the time discretization in the integral operator.
The functional defined in (5.1) provides a framework for analyzing the inverse source problem via an optimization-based approach. Its construction is consistent with the Lipschitz-type stability property established in Appendix A, thereby ensuring the theoretical soundness of the formulation. Thus, the selection of this functional is motivated by the inequality stated in Appendix A, which is associated with the Lipschitz stability result for the inverse source problem presented in Problem 1.
(a) first component
(b) rel. error $6,9\%$
(c) rel. error $43,2\%$
(d) second component
(e) rel. error $14,4\%$
(f) rel. error $14,4\%$
(g) first component
(h) rel. error $4,1\%$
(i) rel. error $63,6\%$
(j) second component
(k) rel. error $7,7\%$
(l) rel. error $7,7\%$
Figure 4. Source reconstruction for $F_{3}(x,y)$ using a coupling matrix given by $q_{11}=1$, $q_{12}=4$, $q_{21}=0$, and $q_{22}=1$. The first column displays the sources and the locations of the observation domains, indicated by a white dashed line. The second column displays the reconstructions from observing both components. The third column shows the reconstruction obtained by observing only the last component.
(a) 1st component
(b) rel. error $17,7\%$
(c) rel. error $20,4\%$
(d) rel. error $36,1\%$
(e) 2nd component
(f) rel. error $14,3\%$
(g) rel. error $27,9\%$
(h) rel. error $22,0\%$
(i) 1st component
(j) rel. error $8,1\%$
(k) rel. error $9,0\%$
(l) rel. error $36,5\%$
(m) 2nd component
(n) rel. error $6,2\%$
(o) rel. error $11,6\%$
(p) rel. error $25,8\%$
Figure 5. Source reconstruction for $F_{3}(x,y)$ using the coupling matrix $q_{11}=q_{22}=0$, $q_{12}=4$ and $q_{21}=2$. The first column displays the sources and the locations of the observation domains, indicated by a white dashed line. The second, third, and fourth columns display the reconstructions with both components, followed by the reconstructions with the first and second components, respectively.
@@ 6. Conclusions
In this paper, we introduced and analyzed inverse problems aimed at recovering spatial source distributions in coupled parabolic systems. We addressed two types of coupling: one involving a potential matrix with constant components and another involving a potential matrix with spatially dependent entries A significant contribution of our work is the development of source reconstruction algorithms based on local internal measurements from one scalar component of the system’s state. To the best of our knowledge, there are no previous studies on reconstruction algorithms for these sources.
One of the primary advantages of our approach is the minimal requirement for the measured state variables needed to reconstruct the source coefficients. Nevertheless, several related theoretical questions remain unresolved, particularly those concerning stability and uniqueness of the inverse source problem using the methodology via spectral analysis, control null results, and Volterra equations. This lack of resolution is not unexpected; even in the scalar case analyzed in [25], achieving stability required stringent assumptions on the temporal derivative $\sigma^{\prime}(t)$. Specifically, a stability result was obtained when $\sigma^{\prime}(t)$ decreases exponentially in $(0,T)$, i.e., $\sigma^{\prime}(t)\approx e^{-1/t}$ (see [25, Proof of Theorem 1.3, Step 4]). This particular choice of $\sigma^{\prime}(t)$ is closely linked to the null controllability problem since the control function cost has an exponential dependence of the form $e^{1/T}$ (see [25, Lemma 1.1]). Consequently, establishing stability via our methodology would similarly require precise control-cost estimates for coupled parabolic systems. Moreover, such estimates must simultaneously allow appropriate bounds on $\sigma(t)$ and $\sigma^{\prime}(t)$ while ensuring $\sigma(T)\neq 0$. Currently, however, obtaining these estimates remains an open problem, as the control function cost involves an exponential term of the form $e^{T+1/T}$. To provide a positive answer regarding the stability property, we have included a Lipschitz-type stability result for Problem 1 via Carleman inequalities in the Appendix A.